# Robust and Adaptive Deep Reinforcement Learning for Enhancing Flow Control around a Square Cylinder with Varying Reynolds Numbers

### v2

@@ I Introduction
Deep learning (DL) has demonstrated remarkable proficiency in extracting nonlinear features from complex systems, such as chaotic systems.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015) DL can learn and represent highly abstract features, thereby capturing the nonlinear patterns and dynamic changes inherent in complex systems.Mahesh (2020); Sarker (2021); Burrell (2016) This feature enables DL to effectively extract features of complex systems.Sarker (2021); Burrell (2016) Reinforcement learning (RL) is a key subfield of machine learning that has attracted significant attention in recent years, particularly after the integration of deep neural networks to create the Deep Q-Network (DQN) artificial intelligence agent. The DQN, by learning from high-dimensional inputs, has demonstrated remarkable performance in various complex tasks, marking an important milestone.Mnih et al. (2015) The triumph of AlphaGo over the world Go champion in 2016 further highlighted RL’s adeptness at complex decision-making and strategic planning.Silver et al. (2017) These developments have significantly propelled the interest and perceived impact of DRL within the AI domain. The synergistic ability of DL combined with RL can solve decision-making problems that were previously considered intractable, especially in non-linear and high-dimensional fusion scenarios.Kaelbling, Littman, and Moore (1996); Arulkumaran et al. (2017)
@@ III.2 Control effect based on SAC algorithm
At higher $Re$, similar learning trajectories are evident in Figs. 6(b), 6(c) and 6(d), where rewards in each instance converge over time. This consistency underscores the remarkable proficiency of DRL agents in acquiring effective control methods across a spectrum of $Re$. Beyond this commonality, it is observable that the extent of variation within each reward trajectory differs. With an increase in the $Re$, the level of fluctuation intensifies. Specifically, at $Re=400$, the variability remains comparatively large at a notable magnitude even while the average reward stabilizes. This extensive fluctuation in the reward signal at higher $Re$ suggests that more complex flow dynamics present greater challenges, necessitating more sophisticated control strategies.
@@ III.3 Detailed flow field comparison before and after control
Figure 11: The instantaneous velocity contour plots from top to bottom correspond to $Re=100$, 200, 300, and 400, respectively. Here, $t_{0}$ represents the initial time instant, depicting the velocity field before control, $t_{1}$ represents the velocity field during the control process, and $t_{2}$ represents the velocity field after the control has ceased, signifying the outcome of flow control based on the DRL algorithm.

### v3

@@ I Introduction
Deep learning (DL) has demonstrated remarkable proficiency in extracting nonlinear features from complex systems, such as chaotic systems.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015) DL can learn and represent highly abstract features, thereby capturing the nonlinear patterns and dynamic changes inherent in complex systems.Batta (2020); Sarker (2021); Burrell (2016) This feature enables DL to effectively extract features of complex systems.Sarker (2021); Burrell (2016) Reinforcement learning (RL) is a key subfield of machine learning that has attracted significant attention in recent years, particularly after the integration of deep neural networks to create the Deep Q-Network (DQN) artificial intelligence agent. The DQN, by learning from high-dimensional inputs, has demonstrated remarkable performance in various complex tasks, marking an important milestone.Mnih et al. (2015) The triumph of AlphaGo over the world Go champion in 2016 further highlighted RL’s adeptness at complex decision-making and strategic planning.Silver et al. (2017) These developments have significantly propelled the interest and perceived impact of DRL within the AI domain. The synergistic ability of DL combined with RL can solve decision-making problems that were previously considered intractable, especially in non-linear and high-dimensional fusion scenarios.Kaelbling, Littman, and Moore (1996); Arulkumaran et al. (2017)
@@ III.2 Control effect based on SAC algorithm
At higher $Re$, similar learning trajectories are evident in Figs. 6(b), 6(c) and 6, where rewards in each instance converge over time. This consistency underscores the remarkable proficiency of DRL agents in acquiring effective control methods across a spectrum of $Re$. Beyond this commonality, it is observable that the extent of variation within each reward trajectory differs. With an increase in the $Re$, the level of fluctuation intensifies. Specifically, at $Re=400$, the variability remains comparatively large at a notable magnitude even while the average reward stabilizes. This extensive fluctuation in the reward signal at higher $Re$ suggests that more complex flow dynamics present greater challenges, necessitating more sophisticated control strategies.
@@ III.3 Detailed flow field comparison before and after control
Figure 11: Instantaneous velocity contours for a range of Reynolds numbers ($Re=100$, 200, 300, and 400) captured at successive temporal instances. Here, $t_{0}$ denotes the initial time, illustrating the uncontrolled velocity field. $t_{1}$ corresponds to the controlled state during the implementation of the DRL-based flow control algorithm. $t_{2}$ conveys the velocity field following the cessation of control, thereby reflecting the results of the DRL algorithm’s application. (a) $Re=100$, at $t_{0}$. (b) $Re=100$, at $t_{1}$. (c) $Re=100$, at $t_{1}$. (d) $Re=200$, at $t_{0}$. (e) $Re=200$, at $t_{1}$. (f) $Re=200$, at $t_{1}$. (g) $Re=300$, at $t_{0}$. (h) $Re=300$, at $t_{1}$. (i) $Re=300$, at $t_{1}$. (j) $Re=400$, at $t_{0}$. (k) $Re=400$, at $t_{1}$. (l) $Re=400$, at $t_{1}$.